# Logarithmic Smoothing for Adaptive PAC-Bayesian Off-Policy Learning

### v1

@@ Abstract
Off-policy learning serves as the primary framework for learning optimal policies from logged interactions collected under a static behavior policy. In this work, we investigate the more practical and flexible setting of adaptive off-policy learning, where policies are iteratively refined and re-deployed to collect higher-quality data. Building on the success of PAC-Bayesian learning with Logarithmic Smoothing (LS) in static settings, we extend this framework to the adaptive scenario using tools from online PAC-Bayesian theory. Furthermore, we demonstrate that a principled adjustment to the LS estimator naturally accommodates multiple rounds of deployment and yields faster convergence rates under mild conditions. Our method matches the performance of leading offline approaches in static settings, and significantly outperforms them when intermediate policy deployments are allowed. Empirical evaluations across diverse scenarios highlight both the advantages of adaptive data collection and the strength of the PAC-Bayesian formulation.
@@ 1 Introduction
Off-policy learning has become a ubiquitous task for practitioners in a wide range of applications such as recommender systems (Sakhi_Rohde_Gilotte_2023, ), ad placement (bottou2013counterfactual, ), and personalized medicine (kallus2018policy, ). This task involves learning optimal policies from offline data summarizing a behavior policy’s past interactions. These interactions, typically captured as tuples of contextual features, action played, and reward received, hold valuable insight into the underlying dynamics of the environment. Each tuple represents a single round of interaction, where the agent observes a context (including relevant features), takes an action according to its current, behavior policy, and receives a reward that depends on both the observed context and the taken action.
@@
The off-policy learning literature focused on developing efficient approaches to learn from these logged interactions bottou2013counterfactual ; swaminathan2015batch ; london2019bayesian ; sakhi2023pac . At the core of these methods the development of tailored off-policy estimators with improved concentration properties (bottou2013counterfactual, ; aouali23a, ; metelli2021subgaussian, ; sakhi2024logarithmic, ), combined with tight generalization bounds allowing efficient learning (swaminathan2015batch, ; sakhi2023pac, ; sakhi2024logarithmic, ). The methodologies employed derive from the pessimism principle jin2021pessimism , which has been demonstrated to be optimal for learning in these scenarios. Particularly, analyzing this learning problem from the PAC-Bayesian learning theory lens london2019bayesian proved efficient, improving the seminal work of (swaminathan2015batch, ). This direction was further explored by sakhi2023pac ; aouali23a ; pac_bayes_unified ; gabbianelli2023importance ; sakhi2024logarithmic , reducing assumptions and improving the pessimistic principle, i.e. controlling the worst-case scenario, with latest developments (sakhi2024logarithmic, ) resulting in state of the art pessimistic approaches.
The success of these pessimistic learning approaches fundamentally depend on the quality of the logged data (jin2021pessimism, ; gabbianelli2023importance, ; sakhi2024logarithmic, ), thus the quality of the behavior policy and its ability to play meaningful actions. In particular, it was demonstrated that off-policy learning approaches converge to the optimal policy with enough samples, with faster convergence when the behavior policy explores well the optimal actions (sakhi2024logarithmic, ). To achieve this convergence, a naive strategy consists in collecting more data with the same logging system. However, this approach is limited as it does not consider the benefits of learning from previous interactions. Then, a better adaptive strategy would use the already collected data to refine the design of the behavior policy, adapting better to the underlying dynamics of the problem, and collecting interaction data of higher quality. More precisely, the adaptive strategy performs a sequence of data collection experiments by determining at each round a new behavior policy using interaction data collected during previous experiments. The obtained policy is then deployed for the next round to collect better quality samples.
This adaptive (or sequential) off-policy learning approach recently sparked interest (adaptive, ; scrm, ). However, existing algorithms are not in line with common practices, either requiring strong assumptions such as uniform coverage adaptive ; scrm , or involving unrealistic data collection strategies and cannot accommodate online updates scrm . In the conventional off-policy learning setting, a prominent route exploits PAC-Bayes learning guedj2019primer ; hellstrom2023generalization ; alquier2024user to reach both theory-driven learning algorithm and convergence guarantees sakhi2024logarithmic . Given the existence of online PAC-Bayes methods via supermartingales haddouche2022online ; haddouche2023pac ; chugg2023unified allowing a continuous data stream, a natural question then arise: can we reach PAC-Bayesian adaptive off-policy algorithms?
Intuitively, an adaptive PAC-Bayesian strategy would yield in better behavior policies at each deployment than sakhi2024logarithmic , increasing the quality of the logged samples, which in turn result in a faster convergence towards the optimal policy. In this work, we formalize this intuition and develop the necessary tools to study this strategy and obtain such learning guarantees.
Contributions. By leveraging the online PAC-Bayesian toolbox, we extend in Sec. 3 the offline pessimistic logarithmic smoothing PAC-Bayesian strategy sakhi2024logarithmic to the adaptive scenario adaptive , allowing the approach to handle sequential interaction data, collected by progressively adaptive behavior policies. This helps us define a first algorithm that already exhibits improved performance compared to the offline approach. These improvements do not require sophisticated data collection strategies nor strong assumptions. Additionally, we identify a core limitation of the logarithmic smoothing estimator and propose in Sec. 4 an adjustment to address it. This correction allows for faster convergence rates to the optimal policy under reasonable and interpretable assumptions. Our proposed approach matches the guarantees of the offline setting and improves upon them when we have the possibility to deploy multiple policies, completely generalizing the state-of-the-art offline PAC-Bayesian approach and allowing for adaptive PAC-Bayesian off-policy learning. We also conduct in Sec. 5 numerical experiments to demonstrate the effectiveness of our approach and quantify how it improves upon the pure offline setting as well as recent adaptive strategies.
@@
Outline. Sec. 2 introduces the adaptive framework and provides the necessary background. In Sec. 3, we provide a first extension to the logarithmic smoothing (LS) PAC-Bayesian off-policy learning approach to the adaptive setting, study its benefits and limitation. In Sec. 4, we provide a simple adjustment to the LS estimator, achieving an accelerated convergence rate under mild conditions. We also study its behavior under different data collection strategies and demonstrate its effectiveness for all scenarios. Empirical validation is conducted in Sec. 5 to highlight the favorable performance of the our proposed method, and Sec. 6 provides concluding remarks.
@@ 2 Framework
Adaptive Contextual bandits.
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the context space, which is a compact subset of $\mathbb{R}^{d}$, and let $\mathcal{A}=[K]$ be a finite action set. An agent’s actions are guided by a stochastic policy $\pi:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{A})$ within a policy space $\Pi$. Given a context $x\in\mathcal{X},\pi(\cdot\mid x)$ is a probability distribution over the action set $\mathcal{A};\pi(a\mid x)$ is the probability that the agent selects action $a$ in context $x$. In a general adaptive framework, we consider that the agent interacts sequentially with a data stream: at time $k\geq 0$, the agent deploys a new behavior policy $\pi_{k}$ and a new batch of contexts $\{x_{j}\mid j\in I_{k}\}$ observed by the agent is revealed, where $|I_{k}|=n_{k}$ and all $(I_{k})_{k\geq 0}$ are disjoint. These contexts are drawn i.i.d. from a fixed, unknown distribution $\nu$. For each $j\in I_{k}$, the agent selects for context $x_{j}$ an action $a_{j}\sim\pi_{k}\left(\cdot\mid x_{j}\right)$, where $\pi_{k}$ is the current behavior policy of the agent at time $k$, and receives a stochastic cost $c_{j}\in[-1,0]$ that depends on the observed context $x_{j}$ and the taken action $a_{j}$. This cost $c_{j}$ is sampled from an unknown cost distribution $p\left(\cdot|x_{j},a_{j}\right)$ and we denote by $c(a,x)$ its expectation under $p\left(\cdot|x_{j},a_{j}\right)$. After $|I_{k}|=n_{k}$ interactions, the agent adaptively updates its behavior to $\pi_{k+1}$ w.r.t. previous behavior policies $(\pi_{j})_{j=0\cdots k}$ and is allowed to use all available data $\{x_{i},a_{i},c_{i}\}_{i=1\cdots N_{k}}$, where $N_{k}:=\sum_{i=0}^{k}n_{i}$. To ease notations, we denote by $P(\pi):=\nu\times\pi\times p$ the distribution of $(x,a,c)$ tuples generated by policy $\pi$. The performance of any policy $\pi$ is assessed by its risk, which is defined as the expectation under the context distribution $\nu$, and the played actions $a$ by $\pi$ of the expected costs $c(x,a)$, as in:
@@
In this work, our goal is to come up with adaptive learning algorithms that guarantee the convergence to the optimal policy $\pi_{\star}$, defined as the policy $\pi$ minimizing the risk as: $\pi_{\star}=\operatorname*{arg\,min}_{\pi\in\Pi}R(\pi)\,.$
PAC-Bayes learning. To learn the optimal policy $\pi_{\star}$, we employ the pessimism principle (jin2021pessimism, ), shown to be optimal for off-policy learning jin2021pessimism . Specifically, we aim at minimizing an empirical, tight upper bound on the risk $R(\pi)$ over policies $\pi\in\Pi$. To do so, we leverage PAC-Bayes learning (alquier2024user, ) which controls the generalization guarantee of a posterior distribution $Q$ over a predictor space $\Theta$ (namely $Q\in\mathcal{P}(\Theta)$). To make this consistent with controlling the risk of a policy $\pi$, we exploit a specific parametrization of policies $\pi\in\Pi$. For any $\theta\in\Theta$, we define the deterministic policy: $\forall x,d_{\theta}(x):=\mathds{1}\{a=f_{\theta}(x)\}$, where $f_{\theta}:\mathcal{X}\rightarrow\mathcal{A}$, a parametrized predictor. Then, each distribution $Q\in\mathcal{P}(\Theta)$ defines a policy $\pi_{Q}$ by setting the probability of action $a$ given context $x$ as the probability that a random predictor $f_{\theta}\sim Q$ maps $x$ to action $a$, that is:
@@
This expression is not restrictive as any policy can be represented in this form (sakhi2023pac, , Theorem 3.1). With this characterization, the space of policies considered $\Pi$ can be entirely described by a predictor space $\Theta$ and a set of distributions over it $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$ sakhi2023pac . With this duality, we can obtain with the PAC-Bayesian toolbox, by fixing a reference distribution $P$, and setting $\delta\in(0,1]$ a bound of the following general form, holding with probability $1-\delta$, simultaneously for all $Q$:
@@
upper bounding the true risk of $\pi_{Q}$ by a a risk estimator $\hat{R}(\pi_{Q})$ and $\operatorname{KL}(Q||P):=\mathbb{E}_{Q}\left[\log\left(\frac{dQ}{dP}\right)\right]$ is the Kullback-Leibler divergence between $P$ and $Q$. The tightness of the bound fundamentally depends on the concentration properties of the chosen estimator $\hat{R}$ sakhi2024logarithmic .
Adaptive algorithms. We define the full (countable) dataset $S:=\cup_{k\in\mathbb{N}}S_{k}$, where $S_{k}:=\left\{(x_{j},a_{j},c_{j})\mid j\in I_{k}\right\}$ is the set of logged data at time $k$. We also define $\mathcal{F}_{k}$ a filtration adapted to $\left(\cup_{j=0}^{k}S_{j}\right)_{j\in 0\cdots k}$ ($\mathcal{F}_{-1}$ is the trivial $\sigma$-algebra). This framework generalizes the pure static case, on which is based sakhi2024logarithmic , as $S_{0}$ summarizes the interactions of a single data batch with the starting policy $\pi_{0}$. We denote as adaptive algorithm, any algorithm Alg such that its output $\pi_{k+1}$ at time $k$ is $\mathcal{F}_{k}$-measurable. In other words, if we denote all the collected interactions at time $k$ as $S_{0:k}:=\cup_{\ell=0}^{k}S_{\ell}$, , we can express $\pi_{k+1}=\texttt{Alg}(S_{0:k})$. We assume that for each $k$, $(x_{j})_{j\in I_{k}}\sim\nu^{\otimes n_{k}}$ is i.i.d. and that, conditionally to $\mathcal{F}_{k-1}$, $S_{k}$ is constituted of mutually independent random variables.
Regularized IPS. Given the successive arrival of data batches $(S_{k})_{k\geq 0}$ and for all $j\geq 0$, $\pi_{j+1}=\texttt{Alg}(S_{0:j})$ is given by an adaptive algorithm, we can leverage the collected data at any time $k\geq 0$ to define an Inverse Propensity Scoring (IPS) estimator horvitz1952generalization , for the risk of any policy $\pi$: $\hat{R}_{0:k}^{\textsc{IPS}}(\pi)=\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}\frac{\pi(a_{i}|x_{i})}{\pi_{j}(a_{i}|x_{i})}c_{i}\,,$ which recovers the static case when $k=0$. Even with the potential dependencies between the behavior policies $\pi_{j}$, this estimator is unbiased under the common support assumption mcbook . However, as in the static case, this estimator has poor concentration properties metelli2021subgaussian and regularized variants, trading some bias for lower variance are preferred bottou2013counterfactual ; optimistic_shrinkage ; aouali23a ; sakhi2024logarithmic . The majority of these regularized estimators can be defined by specific choices of a regularizer $h:[0,1]^{2}\times[-1,0]\rightarrow\mathbb{R}^{-}$ in the following general expression:
$\hat{R}_{0:k}^{h}(\pi)=\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}h(\pi(a_{i}|x_{i}),\pi_{j}(a_{i}|x_{i}),c_{i})\,,$
with $h$ respecting $\forall(p,q,c)\in[0,1]^{2}\times[-1,0]\,,\quad pc/q\leq h(p,q,c)\leq 0$. $h(p,q,c)=pc/q$ recovers the IPS estimator while other choices recover known, well behaved estimators such as clipped IPS bottou2013counterfactual , harmonic metelli2021subgaussian and Implicit-Exploration gabbianelli2023importance to name a few. When $k=0$, we denote $\hat{R}_{0:0}^{h}$ as $\hat{R}_{0}^{h}$.

### v2

# Sequential Off-Policy Learning with Logarithmic Smoothing

@@ Abstract
Off-policy learning enables training policies from logged interaction data. Most prior work considers the batch setting, where a policy is learned from data generated by a single behavior policy. In real systems, however, policies are updated and redeployed repeatedly, each time training on all previously collected data while generating new interactions for future updates. This sequential off-policy learning setting is common in practice but remains largely unexplored theoretically. In this work, we present and study a simple algorithm for sequential off-policy learning, combining Logarithmic Smoothing (LS) estimation with online PAC-Bayesian tools. We further show that a principled adjustment to LS improves performance and accelerates convergence under mild conditions. The algorithms introduced generalize previous work: they match state-of-the-art offline approaches in the batch case and substantially outperform them when policies are updated sequentially. Empirical evaluations highlight both the benefits of the sequential framework and the strength of the proposed algorithms.
@@ 1 INTRODUCTION
Off-policy learning has become a ubiquitous task for practitioners in a wide range of applications such as recommender systems (Sakhi et al., 2023b), ad placement (Bottou et al., 2013), and personalized medicine (Kallus and Zhou, 2018). This task involves learning optimal policies from offline data summarizing a behavior policy’s past interactions. These interactions, typically captured as tuples of contextual features, action played, and reward received, hold valuable insight into the underlying dynamics of the environment. Each tuple represents a single round of interaction, where the agent observes a context (including relevant features), takes an action according to its current, behavior policy, and receives a reward that depends on both the observed context and the taken action.
@@
The off-policy learning literature focused on developing efficient approaches to learn from these logged interactions (Bottou et al., 2013; Swaminathan and Joachims, 2015; London and Sandler, 2019; Sakhi et al., 2023a). At the core of these methods lies the development of tailored off-policy estimators with improved concentration properties (Bottou et al., 2013; Aouali et al., 2023; Metelli et al., 2021; Sakhi et al., 2024), combined with tight generalization bounds allowing efficient learning (Swaminathan and Joachims, 2015; Sakhi et al., 2023a, 2024). The methodologies employed derive from the pessimism principle (Jin et al., 2021), which has been demonstrated to be optimal for learning in these scenarios. Particularly, analyzing this learning problem from the PAC-Bayesian learning theory lens (London and Sandler, 2019) proved effective, improving the seminal work of Swaminathan and Joachims (2015). This direction was further explored by Sakhi et al. (2023a); Aouali et al. (2023, 2024); Gabbianelli et al. (2024); Sakhi et al. (2024), reducing assumptions and improving the pessimistic principle, i.e. controlling the worst-case scenario, with latest developments (Sakhi et al., 2024) resulting in state of the art approaches.
These methods are well understood and have been extensively studied in the one-step, batch setting. In practice, however, practitioners in production environments typically apply them sequentially (Zenati et al., 2023), repeatedly updating and redeploying policies by retraining on the full history of collected data while simultaneously generating new interactions for future updates. Intuitively, sequential off-policy learning is preferable, as it can yield improved policies at each deployment compared to batch approaches. This, in turn, enhances the quality of the logged data and accelerates convergence toward the optimal policy. Although this sequential off-policy learning paradigm is ubiquitous in real-world applications, it remains largely underexplored from a theoretical perspective. Recent work has begun to address this gap (Bibaut et al., 2021; Zenati et al., 2023), but existing algorithms deviate from common practice: they rely on strong uniform coverage assumptions and often require unrealistic data collection strategies (Zenati et al., 2023). In the conventional off-policy learning setting, some of these assumptions are avoided via the use of PAC-Bayes learning (Alquier, 2024) combined with novel off-policy estimators to reach both theory-driven learning algorithms and convergence guarantees (Sakhi et al., 2024). However, only batch algorithms have been developed, leaving the existence of sequential off-policy PAC-Bayes algorithms as an open question. In this work, we close this gap and develop the first sequential, off-policy PAC-Bayes algorithms with associated theoretical guarantees.
Contributions. We study the sequential deployment of state-of-the-art off-policy learning methods, a paradigm extensively used in real-world applications (Bibaut et al., 2021; Zenati et al., 2023). We extend in Sec.˜3 the off-policy, logarithmic smoothing PAC-Bayesian strategy of Sakhi et al. (2024) to the sequential scenario, allowing the approach to handle successive interaction data, and progressively update the policy. This sequential strategy, detailed in Algorithm˜1, is theoretically analyzed and shown to benefit from improved performance compared to its offline counterpart. Building on this, we identify a core limitation of our previous algorithm, inherited from Sakhi et al. (2024). In Sec.˜4, we address this issue by introducing a second original sequential algorithm (Algorithm˜2) enjoying accelerated convergence rates to the optimal policy under reasonable assumptions. Our sequential algorithms encompass the offline scenario and strictly generalize the state-of-the-art offline PAC-Bayesian approach. Finally, we conduct in Sec.˜5 numerical experiments to demonstrate the effectiveness of our approach, showing that they outperform both offline strategies and recent sequential strategies.
@@
Outline. Sec.˜2 introduces the sequential framework and the necessary background. Sec.˜3 provably extends the logarithmic smoothing (LS) PAC-Bayesian off-policy learning approach to the sequential setting and studies its benefits and limitations. Sec.˜4 provides a simple adjustment to the LS estimator, achieving an accelerated convergence rate under mild conditions. Empirical validation is conducted in Sec.˜5 to highlight the favorable performance of our proposed method, and Sec.˜6 provides concluding remarks.
@@ 2 FRAMEWORK
Contextual bandits.
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the context space, which is a compact subset of $\mathbb{R}^{d}$, and let $\mathcal{A}=[K]$ be a finite action set. An agent’s actions are guided by a stochastic policy $\pi:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{A})$ within a policy space $\Pi$. Given a context $x\in\mathcal{X},\pi(\cdot\mid x)$ is a probability distribution over the action set $\mathcal{A};\pi(a\mid x)$ is the probability that the agent selects action $a$ in context $x$.
Sequential framework. We consider that the agent interacts sequentially with a data stream: at time $k\geq 0$, the agent deploys a new behavior policy $\pi_{k}$ and a new batch of contexts $\{x_{j}\mid j\in I_{k}\}$ observed by the agent is revealed, where $|I_{k}|=n_{k}$ and all $(I_{k})_{k\geq 0}$ are disjoint. These contexts are drawn i.i.d. from the same, unknown distribution $\nu$.
For each $j\in I_{k}$, the agent selects for context $x_{j}$ an action $a_{j}\sim\pi_{k}\left(\cdot\mid x_{j}\right)$, where $\pi_{k}$ is the current behavior policy of the agent at time $k$, and receives a stochastic cost $c_{j}\in[-1,0]$ that depends on the observed context $x_{j}$ and the taken action $a_{j}$. This cost $c_{j}$ is sampled from an unknown cost distribution $p\left(\cdot|x_{j},a_{j}\right)$ and we denote by $c(a,x)$ its expectation under $p\left(\cdot|x_{j},a_{j}\right)$. After $|I_{k}|=n_{k}$ interactions, the agent sequentially updates its behavior to $\pi_{k+1}$ w.r.t. previous behavior policies $(\pi_{j})_{j=0\cdots k}$ and is allowed to use all available data $\{x_{i},a_{i},c_{i}\}_{i=1\cdots N_{k}}$, where $N_{k}:=\sum_{i=0}^{k}n_{i}$.
Formally, we define the full (countable) dataset $S:=\cup_{k\in\mathbb{N}}S_{k}$, where $S_{k}:=\left\{(x_{j},a_{j},c_{j})\mid j\in I_{k}\right\}$ is the set of logged data at time $k$. We also define $\mathcal{F}_{k}$ a filtration adapted to $\left(\cup_{j=0}^{k}S_{j}\right)_{j\in 0\cdots k}$ ($\mathcal{F}_{-1}$ is the trivial $\sigma$-algebra). We assume that, conditionally to $\mathcal{F}_{k-1}$, $S_{k}$ is constituted of mutually independent random variables. This generalizes the pure static case, corresponding to $k=0$. We denote as sequential algorithm, any algorithm Alg such that its output $\pi_{k+1}$ at time $k$ is $\mathcal{F}_{k}$-measurable. In other words, if we denote $S_{0:k}:=\cup_{j=0}^{k}S_{j}$, then we have $\pi_{k+1}=\texttt{Alg}(S_{0:k})$.
To ease notations, we denote by $P(\pi):=\nu\times\pi\times p$ the distribution of $(x,a,c)$ tuples generated by policy $\pi$. The performance of any policy $\pi$ is assessed by its risk, which is defined as the averaged incurred cost of $\pi$ over the draw of new context:
@@
Our goal is to come up with sequential learning algorithms that guarantee the convergence to the optimal policy $\pi_{\star}$, defined as: $\pi_{\star}=\operatorname*{arg\,min}_{\pi\in\Pi}R(\pi)\,.$ For any context $x$, the optimal action is defined as $a_{\star}(x)=\operatorname*{arg\,min}_{a^{\prime}\in\mathcal{A}}c(a^{\prime},x)$ with ties broken arbitrarily. Hence, the optimal policy $\pi_{\star}$ chooses deterministically the optimal action $a_{\star}(x)$ for every context $x$:
PAC-Bayesian off-policy learning. State of the art off-policy learning algorithms (Sakhi et al., 2023a; Gabbianelli et al., 2024; Sakhi et al., 2024) employ the pessimism principle, shown to be optimal in the batch, off-policy learning setting (Jin et al., 2021). These methods aim at minimizing an empirical, tight upper bound on the risk $R(\pi)$ over policies $\pi\in\Pi$. These upper bounds leverage PAC-Bayes learning (Alquier, 2024) which controls the generalization guarantee of a posterior distribution $Q$ over a predictor space $\Theta$ (namely $Q\in\mathcal{P}(\Theta)$). To make this consistent with controlling the risk of a policy $\pi$, we exploit a specific parametrization of policies $\pi\in\Pi$. For any $\theta\in\Theta$, we define the deterministic policy: $\forall x,d_{\theta}(x):=\mathds{1}\{a=f_{\theta}(x)\}$, where $f_{\theta}:\mathcal{X}\rightarrow\mathcal{A}$, a parametrized predictor. Then, each distribution $Q\in\mathcal{P}(\Theta)$ defines a policy $\pi_{Q}$ by the following mapping:
@@
This expression is general as any policy can be represented in this form (Sakhi et al., 2023a, Theorem 3.1). With this duality, Sakhi et al. (2024) gets with the PAC-Bayesian toolbox, by fixing a reference distribution $P$, and setting $\delta\in(0,1]$ a learning bound of the following general form, holding with probability $1-\delta$, simultaneously for all $Q$:
@@
upper bounding the true risk of $\pi_{Q}$ by a risk estimator $\hat{R}(\pi_{Q})$ and $\operatorname{KL}(Q||P):=\mathbb{E}_{Q}\left[\log\left(\frac{dQ}{dP}\right)\right]$, the Kullback-Leibler divergence between $P$ and $Q$. In the batch setting, these upper bounds are used as learning objectives and the convergence guarantees of their learned policies are studied (Gabbianelli et al., 2024; Sakhi et al., 2023a). We extend this paradigm to the sequential framework, and begin by defining off-policy estimators using the interaction data collected.
Regularized IPS. Given the successive arrival of data batches $(S_{k})_{k\geq 0}$ and output of sequential algorithm $(\pi_{k})_{k\geq 0}$, we define for any $\pi\in\Pi$ the Inverse Propensity Scoring (IPS) estimator (Horvitz and Thompson, 1952) as: $\hat{R}_{0:k}^{\textsc{IPS}}(\pi)=\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}\frac{\pi(a_{i}|x_{i})}{\pi_{j}(a_{i}|x_{i})}c_{i}\,$, encompassing the static case ($k=0$). This estimator is unbiased under the common support assumption (Owen, 2013). However, as in the static case, this estimator has poor concentration properties (Metelli et al., 2021) and regularized variants, trading some bias for lower variance are preferred (Bottou et al., 2013; Su et al., 2020). The majority of these regularized estimators stems from specific choices of a regularizer $h:[0,1]^{2}\times[-1,0]\rightarrow\mathbb{R}^{-}$ in the following definition:
$\hat{R}_{0:k}^{h}(\pi):=\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}h(\pi(a_{i}|x_{i}),\pi_{j}(a_{i}|x_{i}),c_{i})\,$,
with $h$ respecting $\forall(p,q,c)\in[0,1]^{2}\times[-1,0]\,,\quad pc/q\leq h(p,q,c)\leq 0$. $h(p,q,c)=pc/q$ recovers the IPS estimator while other choices recover known, well behaved estimators such as clipped IPS (Bottou et al., 2013), power-mean estimators (Metelli et al., 2021) and Implicit-Exploration (Gabbianelli et al., 2024) to name a few. When $k=0$, we denote $\hat{R}_{0:0}^{h}$ as $\hat{R}_{0}^{h}$.